OpenStudy (aravindg):

3. A block of mass M is pulled along a horizontal frictionless by a rope of mass m. If a force P is applied at the free end of the rope, what will be the force exerted by the rope on the block?

OpenStudy (jamesj):

What do you think it is?

OpenStudy (aravindg):

i dunno

OpenStudy (jamesj):

Every force has an equal and opposite reaction.

OpenStudy (aravindg):

soo

OpenStudy (aravindg):

so

OpenStudy (aravindg):

noo

OpenStudy (jamesj):

Actually we want this: P = (M+m)a, hence a = P/(M+m) thus the force on the block is Ma = MP/(M+m)

OpenStudy (aravindg):

ya thts the one